# Puffy Venuses: the Mass - Radius Impact of Carbon-Rich Atmospheres on Lava Worlds

### v1

@@ 4.1 The puffy Venus case for TOI-561 b
As shown in section 3, both a greater bulk volatile abundance and a higher C/O (or a reducing mantle) lead to more radius inflation. This introduces a degeneracy: the M-R measurements of 55 Cnc e can be explained by both a lower EH abundance with a reducing mantle, or a 1 wtC + 1 wtH abundance with an oxidizing mantle. The case offering the best fit across different C/O seems to be a planet with 1 wtC and no H. Future atmospheric C/O constraints can thus break this degeneracy and offer insight into both the bulk volatile inventory and the mantle redox of this world. Finally, we note that a CI carbon abundance provides more than enough puffiness for 55 Cnc e regardless of C/O. In other words, invoking super-chondritic C abundance is unnecessary in accounting for 55 Cnc e’s radius inflation.
@@ 5 Summary and Conclusion
We present a model for a theoretical class of exoplanets - puffy Venuses - characterized by thick, carbon-rich atmospheres and global magma oceans. The model generates mass-radius relations of puffy Venuses while accounting for C-H-O thermochemistry in the atmosphere and volatile dissolution in the magma ocean. Here are our main findings -
@@
Our findings highlight the observational significance of puffy Venuses in two ways. Firstly, this interior scenario can explain the abnormal puffiness of some rocky exoplanets with chondritic rather than icy or exotic compositions. Conversely, a puffy Venus interpretation of 55 Cnc e reminds us that rocky exoplanets could host $\sim$2 orders of magnitude more carbon than the Earth; assuming Earth-like volatile content for them is inadequate. Secondly, their atmospheric composition, in direct equilibrium with a molten mantle, offers insight into their interior. These include their bulk volatile composition and mantle redox state. Ongoing JWST observations of puffy Venus candidates can, therefore, shed light on both aspects and will improve our understanding of the rocky exoplanets’ formation and evolution.

### v2

@@ 4.1 The puffy Venus case for TOI-561 b
As shown in section 3, both a greater bulk volatile abundance and a higher C/O (or a reducing mantle) lead to more radius inflation. This introduces a degeneracy: the M-R measurements of 55 Cnc e can be explained by both a lower EC abundance with a reducing mantle, or a 1 wtC + 1 wtH abundance with an oxidizing mantle. The case offering the best fit across different C/O seems to be a planet with 1 wtC and no H. Future atmospheric C/O constraints can thus break this degeneracy and offer insight into both the bulk volatile inventory and the mantle redox of this world. Finally, we note that a CI carbon abundance provides more than enough puffiness for 55 Cnc e regardless of C/O. In other words, invoking super-chondritic C abundance is unnecessary in accounting for 55 Cnc e’s radius inflation.
4.3 Caveats and Future Directions
We have made simplifying assumptions in our model in order to make our approach tractable while preserving its generality, especially in parameter spaces with scant experimental data. In this section, we explain our choices and their corresponding caveats.
4.3.1 Opacities
For our semi-grey atmosphere, we choose to use constant opacities $\kappa_{th}$ and $\kappa_{v}$ benchmarked to the Venusian profile. This approach is chosen over a more complex full-physics treatment for necessity and simplicity. Our work is not focused on predicting detailed spectroscopic observations but rather on the lower-order task of constraining maximum atmospheric heights. Thus, this treatment allows us to draw clear population-level insights.
In realistic atmospheres, $\kappa_{th}$ and $\gamma$ are influenced by atmospheric composition, the stellar radiation spectrum, and more complex effects such as clouds and photochemical hazes. While capturing these nuances lies outside the scope of this work, we tested the range of $\kappa_{th}$ and $\gamma$ where our main conclusion qualitatively applies (see fig. 5, right). We calculate Earth-mass planets of fiducial $T_{eq}$ and volatile abundance, with $\kappa_{th}$ down to 1% fiducial and $\gamma$ up to 60×$\times\gamma_{0}$. We compare the radius inflation of these planets with the maximum radius inflation that iron depletion can cause to a rocky world, or the rocky threshold radius (RTR, black dashed line in fig. 5, right). Our atmospheres producing greater radius inflation than the RTR would support our thesis that puffy Venuses have an M-R qualitatively distinct from those of terrestrial planets. We find that an atmosphere with $0.05\times\kappa_{th}$ can still produce planets less dense than RTR with a wide range of $\gamma$. Increasing $\gamma$ leads to decreasing $z_{atm}$ for $\gamma\lesssim 40$. Yet at higher $\gamma$ this trend reverses. This is due to the surface temperature becoming low enough that further increasing $\gamma$ results in shallower magma oceans, forcing more volatiles to the atmosphere and increasing $z_{atm}$. Therefore, our main conclusion is robust to more than 1 order of magnitude of change in $\kappa_{th}$ and $\gamma$.
4.3.2 Refractory C reservoirs
$\mathrm{C(graphite)+O_{2}(gas)\Leftrightarrow CO_{2}(gas)}.$ (46)
The associated equilibrium constant is:
$K_{eq,CCO}=\frac{f_{\mathrm{CO_{2}}}}{a_{C}f_{\mathrm{O_{2}}}}.$ (47)
$K_{eq,CCO}$ rises sharply with temperature; therefore, at a higher T, a greater $f_{\mathrm{CO_{2}}}$ is permitted before reaching the graphite saturation. We calculate $K_{eq,CCO}(T)$ throughout the atmosphere using the tabulated JANAF data(Chase, 1998) for graphite, and calculate $a_{C}$ via the above relation. We find $a_{C}\ll 1$ for most of our planets (see fig. 6). For instance, for planets with OC abundance at $T_{eq}$ = 1500 K, $a_{C}<0.003$, while for dry CI abundance at 2000 K, $a_{C}<0.026$.
Only in the most C-rich (CI abundances) and cool ($T_{eq}\leq$ 1500 K) worlds do we expect graphite saturation, and only for high mass planets (fig. 6, middle). For the dry CI case at $T_{eq}$ = 1000 K (table 1), graphite saturation is present in $>8.4M_{\oplus}$ planets, while for the wet CI, 1000 K case, this is reached for $>2M_{\oplus}$ planets. For the wet CI, 1500 K case, graphite saturation is reached for $>12M_{\oplus}$ worlds. Meanwhile, most radius inflation occurs on hot, low-mass planets. Therefore, graphite precipitation in the atmosphere or near the MO surface is not a concern for the planets we examine.
As Hirschmann (2012) showed, increasing pressure decreases the threshold in carbon abundance for graphite saturation in the mantle. Therefore, our CCO calculation here does not preclude the possibility of graphite/diamond precipitation at great depths. However, the decrease in the graphite saturation threshold is gradual - from 0 to 24 GPa, it decreased by 3$\sim$ 3$\times$ 3 ×. Therefore, if carbon burial through graphite saturation does occur, it likely only affects the cold, highly reduced planets, while the hotter planets ($T_{eq}\geq$ 1500 K), where we predict the most radius inflation, are likely spared of these effects.
[…]
4.3.3 Mantle thermal model
4.3.4 Thermochemistry at the MO - Atmosphere interface
The high P-T environment near the MO surface would likely facilitate more complex thermochemistry than our simplified reaction network permits, such as the further thermal dissociation of molecules to ionic and atomic species like OH-, C, H, and O, or the polymerization of CO and CO2 (Li et al., 2021). While we recognize that previous semi-analytical calculations for exoplanet atmospheric chemistry have covered more species (e.g. Heng & Tsai 2016; Tian & Heng 2024), we choose this simplified reaction network for two reasons. On the one hand, we are limited by the experimentally determined solubility relations. These relations are developed with the mantle outgassing of the Earth in mind. Thus, they span a limited number of volatile species (see, for instance, Liu et al. (2005); Hirschmann et al. (2012); Yoshioka et al. (2019)).
On the other hand, our reaction network is both easily interpretable and complex enough to capture the salient features of $\mathrm{CO_{2}}$ and $\mathrm{H_{2}O}$ thermal dissociation at around 4000 K. A similar argument for a balance between transparency and nuance was used to justify similar reaction networks in Schlichting & Young (2022); Young et al. (2023).
A more extensive reaction network can alter the P-T profiles of the atmosphere by impacting its compressibility and adiabatic lapse rate, as well as the overall volatile partitioning behavior between the two reservoirs. If the net effect of the further thermal dissociation is analogous to the dissociation of CO2 and H2O into CO, O2, and H2, which our model captures, then the atmospheric heights are underestimated by our model. However, more experimental work on the partitioning behaviors of these ionic and atomic species is needed.
More fundamentally, future work on magma ocean planets should move beyond the assumption of a well-defined atmosphere-magma ocean boundary. Recent works on magma ocean sub-Neptunes suggest that boundaries are blurred in two ways. Firstly, rock vapor may be chemically stable near the base of the atmosphere, where the temperature exceeds 5000 K (Misener & Schlichting, 2022, 2023) - a condition produced by our model for $T_{eq}\geq$ 1500 K. Since rock vapor is condensable, its presence in the lower atmosphere could cause a compositional gradient and impact its PT structure (Misener & Schlichting, 2023). Secondly, the MO surface conditions on the hottest planets we probed exceed the critical point for Earth’s mantle material, 80 to 130 MPa and 6500 to 7000 K (Caracas & Stewart, 2023). A puffy, supercritical upper mantle can lower the bulk planet density and potentially alter the mixing behavior of the “dissolved” volatiles. Ultimately, the exotic conditions at the MO surface, where reactive volatiles coexist with supercritical silicates, call for further ab initio simulations and experimental studies.
4.3.5 High atmosphere conditions
@@ 5 Summary and Conclusion
We present a model for a theoretical class of exoplanets - puffy Venuses - characterized by thick, carbon-rich atmospheres and global magma oceans. We advocate this Scenario as a promising explanation for underdense lava worlds. Our model generates mass-radius relations of puffy Venuses while accounting for C-H-O thermochemistry in the atmosphere and volatile dissolution in the magma ocean. Here are our main findings -
@@
Our findings highlight the observational significance of puffy Venuses in two ways. Firstly, this interior Scenario can explain the abnormally large radius of some low-mass exoplanets with chondritic rather than icy or exotic compositions. Conversely, a puffy Venus interpretation of 55 Cnc e reminds us that rocky exoplanets could host $\sim$2 orders of magnitude more carbon than the Earth; assuming Earth-like volatile content for them is inadequate. Secondly, their atmospheric composition, in direct equilibrium with a molten mantle, offers insight into their interior. These include their bulk volatile composition and mantle redox state. Ongoing JWST observations of puffy Venus candidates can, therefore, shed light on both aspects and will improve our understanding of the rocky exoplanets’ formation and evolution.